OpenStudy (anonymous):

Hey guys Does anyone can help me with this? H9P1: RESPONSE TO A DELAYED IMPULSE The current source puts out an impulse of area A=2/π=0.64 Coulombs at time t=9.0s. At t=0 the state is: vC(0)=0.0 and iL(0)=1.0. The equation governing the evolution of the inductor current in this circuit is d2iL(t)dt2+1LCiL(t)=ALCδ(t−9.0) and this H9P2 SOC Consider the second order low pass filter circuit shown below.?

OpenStudy (anonymous):

about H9P2.. if anyone can tell me the formula. I'll be thankful and H9P1 how can I solve it ?

OpenStudy (anonymous):

H9P1 1 0.25 2 22.5 3 22.5 4 5.51091059616e-16 5 -70.6858347058 6 0 7 0 8 0 H9P2 1 L*C 2 L/R 3 1 4 sqrt(1/(L*C))/(2*pi) 5 1/(2*R*C) 6 (1/R)*sqrt(L/C) 7 2*sqrt(L/C)

OpenStudy (anonymous):

CHEK BEST RESPONSE IF U GET RIGHT ANSWERS.

OpenStudy (anonymous):

(1/R)*sqrt(L/C) 2*sqrt(L/C) these are wrong answers, plz let us know the correct one acid16

OpenStudy (anonymous):

thats answers is working for me... i dont know what is wrong... check them again..

OpenStudy (anonymous):

I ve checked 4 to 5 times but it doesn't work........ probably a changed question! of what is it the formula 2*sqrt(L/C)?

OpenStudy (anonymous):

And what's the formula to calculate H9P1 2ND QUESTION? from which u got the following ans... 2... 22.5 can you plz let me know acid16

OpenStudy (anonymous):

give me a sec..

OpenStudy (anonymous):

the answer 22.5 both of them and the answer -70.6858347058 of H9P1 are not correct here .. idk why... and H9P2 (1/R)*sqrt(L/C) 2*sqrt(L/C) are not correct too.. do you know why ?

OpenStudy (anonymous):

H9P1 1 sqrt(1/(L*C)) / (2*pi) 2 ½ * L * iL^2 3 the same as 2 4 iL (t) = cos(2*pi*f*t) 5 vC(t) = -L*2*pi*f*sin(2*pi*f*t)

OpenStudy (anonymous):

just try this formulas, maybe u have diferent values..

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

thanks alot!

OpenStudy (anonymous):

I ve checked 4 to 5 times but it doesn't work........ probably a changed question! of what is it the formula 2*sqrt(L/C) acid16?

OpenStudy (anonymous):

yeeah.. it's not this answer !

OpenStudy (anonymous):

I'm having problems with 2 22.5 3 22.5 5 -70.6858347058

OpenStudy (anonymous):

and the formulas (1/R)*sqrt(L/C) 2*sqrt(L/C).. it doesn't work n' idk why

OpenStudy (anonymous):

h9p1 is solved but P2 is not having green tick for the last 2 acid16 thanx alot

OpenStudy (anonymous):

the formulas are R*sqrt(C/L) 0.5*sqrt(L/C)

OpenStudy (anonymous):

thank you so much !